Which gas phase reaction could probably not be stu Explain too please
Answer #1

answer is (d) since both sides have same number of gaseous moles, there will be no change in the jump of pressure

P1 V1 / T1 = P2 V2 /T2

V2 = 3V1

T2 =2T1

P1 V1 /T1 = P2 x 3V1 / 2T1

P2 = 2/3 P1

OPTION-d